Find r.ns where s is the surface of the sphere x^2+y^2+z^2=a
1
Expert's answer
2020-12-03T19:14:57-0500
"Solution"

To find r.ns where s is the surface of the sphere "x^2+y^2+z^2=a"

We know that "x^2+y^2+z^2=a^2"

Let "a=1" because "a^2=a=1"

We will compute the surface integral "\intop \intop_s x^2 \ ds" where s is the unit sphere.


Apply the parametric representation


"x=sin\ \theta\ cos\ \phi;\ y=sin\ \theta\ sin\ \phi;\ z=cos\ \phi\\
0 \le \phi \le\pi,\ 0 \le \theta \le2\pi,\"

That is

"r(\phi, \theta)=sin\ \phi\ cos\ \theta\ i+sin\ \phi\ sin\ \theta\ j+cos\ \phi\ k"

We then compute

"|r_{\phi} \times r_{\theta}|=sin\ \theta\\
\therefore \intop \intop_Sx^2 dS\\
\implies \intop \intop_D(sin \phi cos \theta)^2|r_{\phi} \times r_{\theta}|dA\\
\implies \intop_0^{2 \pi} \intop_0^{\pi}(sin^2 \phi cos^2 \theta sin \phi)d \phi d\theta\\
\implies \intop_0^{2 \pi}cos^2 \theta d\theta \intop_0^{\pi}(sin^3 \phi )d \phi\\
\implies \intop_0^{2 \pi}\frac12(1+cos2 \theta) d\theta \intop_0^{\pi}(sin \phi-sin \phi\ cos^2 \phi )d \phi\\
\implies \frac12[\theta+\frac12 sin \theta]_0^{2\pi}[-cos \phi + \frac13cos \phi]_0^{\pi}\\
\therefore \frac{4 \pi}{3}"